OpenStudy (anonymous):

Given JKLMN congruent to RSTUV, find MN if NJ=5x-10, UV=6x+3, and VR=3x-6.

OpenStudy (anonymous):

Is there a graph/diagram or something to go along with the question?

OpenStudy (anonymous):

No

OpenStudy (anonymous):

I drew the graph for you. Note that congruent means the two pentagons are exactly the same shape and size. This means that side MN is the same length as UV. And NJ is the same length as VR.

OpenStudy (anonymous):

Would that make MN be 6x+3 ?

OpenStudy (anonymous):

Solve for x first.

OpenStudy (anonymous):

How?

OpenStudy (anonymous):

Equate one side of the first pentagon to the equivalent side of the other pentagon. Then solve for x.

OpenStudy (anonymous):

6x+3=6x+3?

OpenStudy (anonymous):

MN= 6x+3 However you don't know what MN is yet. What other two sides can you make equal to each other? Remember they are the same length.

OpenStudy (anonymous):

5x-10=6x+3=3x-6 ?

OpenStudy (anonymous):

Close but you have taken three sides. Note this pentagon may not be a regular pentagon. Meaning not all of the sides are of equal length.

OpenStudy (anonymous):

6x+3=3x-6?

OpenStudy (anonymous):

The side in one pentagon has to be in the exact same position on the other pentagon. Again you are close, but those two sides are not in the exact same position

OpenStudy (anonymous):

5x-10=3x-6

OpenStudy (anonymous):

Bingo! Now solve for x.

OpenStudy (anonymous):

x=2

OpenStudy (anonymous):

Good now substitute the value for x into the other equation.

OpenStudy (anonymous):

15

OpenStudy (anonymous):

There you go. MN=15

OpenStudy (anonymous):

Thank You

OpenStudy (anonymous):

yw